OpenStudy (anonymous):

Someone please help me figure this out (5th grade math) write words to match the expression 3 + (4x12) 36/4 24 - (6+3) (it's not a multiple choice question) every problem is different I just need help writing words to match the expression.

OpenStudy (anonymous):

@dan815 @darrbreezy

OpenStudy (anonymous):

@darrbreezy Thank you so much can you please help me with another one?

OpenStudy (anonymous):

Ryan has nine 14- ounce bags of popcorn to repackage and sell at the school fair. A small bag holds 3 ounces. How many small bags can he make? (explain the work please) Bianca is making scarves to sell. She has 33 peices of blue fabric, 37 pieces of green fabric, and 41 pieces of red fabric. Suppose Bianca uses 3 pieces of fabric to make 1 scarf. How many scarves can she make/

OpenStudy (triciaal):

help writing words to match the expression example three added to four multiplied by twelve is fifty-one thirty-six divided by four is nine

OpenStudy (triciaal):

example 24 - (6+3) 24 less the sum of nine and three

OpenStudy (anonymous):

@triciaal wouldn't "john had three aples he found four more times twelve" makes sense? I wrote that for the first problem. I think they meant to put a sentence when they said "write words to match the expression"

OpenStudy (triciaal):

nine 14- ounce bags of popcorn to repackage = total available one bag = 3 ozs so how many times can you get 3 ozs from the total you have to repackage?

OpenStudy (anonymous):

42 small bags?

OpenStudy (anonymous):

@darrbreezy No lol I meant for the first question.

OpenStudy (anonymous):

The first word problem.

OpenStudy (triciaal):

@2dreams_ you made up a word problem to match the computation."he found four more times twelve" does not make sense!!

OpenStudy (anonymous):

Oh sorry. I'm kind of confused.

OpenStudy (triciaal):

how about 4 boys gave each of their 12 friends a marble and had three extra, how many marbles did they have altogether

OpenStudy (anonymous):

Is that for the very first problem?

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

so my answer doesn't makes sense at all? sorry i'm kind of struggling with this since yesterday. I'm really really confused.

OpenStudy (triciaal):

For the popcorn it would be 9*14/3 = 3*14 = 42 bags correct

OpenStudy (anonymous):

Ok thank you so much ! Can you help me out with the 3rd one, and if you don't mind I have 2 more.

OpenStudy (triciaal):

scarf (33 + 37 + 41)/3 = total pieces/3 pieces for each 37 scarves

OpenStudy (anonymous):

Jasmine has 8 packs of candle wax to make scented candles. Each pack contains 14 ounces of wax. Jasmine uses 7 ounces of wax to make once candle. How many candles can she make?

OpenStudy (triciaal):

1 pack = 14 ozs 8 * 1 pack = 8 * 14 ozs the amount of ozs available to make candles

OpenStudy (triciaal):

how many ozs do you have to make candles?

OpenStudy (anonymous):

(Last one I promise sorry i kept you here for long.) Maurice puts 130 trading cards in protector sheets he fills 7 sheets and puts the remaining 4 cards in an eighth sheet. Each of the filled sheets has the same number of cards. How many cards are in each filled sheet?

OpenStudy (anonymous):

so you divide?

OpenStudy (triciaal):

I want you to slow down and actually read my posts so you can know how to do the problems.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

14 ozs?

OpenStudy (triciaal):

no read again

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

1 pack has 14 ozs and you have 8 packs so you have 8 times 1 pack so you have 8 times 14 ozs do you see that now?

OpenStudy (anonymous):

yes ma'm

OpenStudy (triciaal):

so what is 8* 14

OpenStudy (anonymous):

112

OpenStudy (triciaal):

so you have 112 ozs of wax and you need 7 ozs of wax to make 1 candle. how many candles can you make?

OpenStudy (triciaal):

in other words, how many times can you get 7ozs out of 112 ozs?

OpenStudy (anonymous):

16?

OpenStudy (triciaal):

16 no ? you need to build your confidence. you can do these. you just need to understand what you need to do and then do.

OpenStudy (anonymous):

do you divide 112 by 7ozs?

OpenStudy (triciaal):

now try the last one

OpenStudy (triciaal):

wait how did you get 16 then if you did not know to do that based on what i explained ?

OpenStudy (anonymous):

you said how many times can i get 7 out of 112 is that what you meant?

OpenStudy (triciaal):

so you understand when I explain it ok

OpenStudy (anonymous):

yes ma'm

OpenStudy (triciaal):

ok now try the last one

OpenStudy (anonymous):

so the answer is 16?

OpenStudy (triciaal):

yes the answer is 16 112 divided by 7 = 16

OpenStudy (anonymous):

You said no

OpenStudy (triciaal):

no I did not!! I said don't use the question mark like you are asking like you are not sure. build your confidence and make a statement.

OpenStudy (anonymous):

Oh ! ok sorry. I thought i was incorrect. I'm really bad at doing math.

OpenStudy (anonymous):

But i"ll try to.

OpenStudy (triciaal):

repost from above 16 no ?

OpenStudy (anonymous):

ok i got you

OpenStudy (triciaal):

that's why it;s important to understand the process and know what to do. if you just throw numbers around anywhere then everything can become chaotic especially when you have to do higher level math.

OpenStudy (triciaal):

now please try the last one i need to get off but i'll finish this with you first

OpenStudy (anonymous):

wait

OpenStudy (triciaal):

Maurice puts 130 trading cards in protector sheets he fills 7 sheets and puts the remaining 4 cards in an eighth sheet. Each of the filled sheets has the same number of cards. How many cards are in each filled sheet?

OpenStudy (anonymous):

so the one before it is 37 right? and sorry last question

OpenStudy (triciaal):

what do you think the question is really asking?

OpenStudy (anonymous):

how many scarves can she makes?

OpenStudy (triciaal):

REPOST scarf (33 + 37 + 41)/3 = total pieces/3 pieces for each 37 scarves

OpenStudy (triciaal):

Let's do the very last one and then you can go back and read everything. I don't appreciate when I take the time to help and the person does not even care to read the post.

OpenStudy (triciaal):

Maurice puts 130 trading cards in protector sheets he fills 7 sheets and puts the remaining 4 cards in an eighth sheet. Each of the filled sheets has the same number of cards. How many cards are in each filled sheet?

OpenStudy (anonymous):

I do care. I'm just really new to this. My teacher wont' explain it at all she just sits on her desk, and gives us worksheets.

OpenStudy (triciaal):

let x be the number of cards on each of the 7 sheets and there are 4 cards on the 8th sheet. He has 130 cards in all 7*x + 4 = 130 to solve for x first isolate the x to one side so take away 4 from each side 7*x + 4 -4 = 130 -4 7*x = 126 you have 7*x to get 1*x divide by 7 7*x/7 = 126/7 same to each side to keep the balance x = 18 this means that each of the 7 sheets had 18 cards each and the 8th sheet had 4 cards for a total of 130 cards. (7 * 18 ) + 4 = 130 read and reread until you get it. I have to leave.

OpenStudy (anonymous):

Thank you.